what are the steps to solve this problem: the length of a rectangular deck is 2 more than 3 times the width. the perimeter is 44 inches,what is the width of the deck?
Answer by checkley77(12844)

L=3W+2
2L+2W=44
2(3W+2)+2W=44
6W+4+2W=44
8W=44-4
8W=40
W=40/8
W=5 ANS. FOR THE WIDTH.
L=3*5+2=15+2=17 ANS. FOR THE LENGTH.
PROOF:
2*17+2*5=44
34+10=44
44=44